40.0 grams of Calcium Carbonate compound (CaCO3) with molar mass of 100g/mole, in 

order to make an aqueous solution that has a mole fraction of solute of 0.200? How much

H2O is needed to achieved the desired result?


1
Expert's answer
2021-03-22T06:14:05-0400

Mass of CaCO3 = 40 g

Molar mass of CaCO3 = 100 g/mol

Moles of CaCO3(n1) = "\dfrac{40}{100}=0.4\space mol"

Mole fraction of solute CaCO3(x1) = 0.200

Moles of solvent H2O = n2

"\dfrac{n_1}{n_1+n_2}=0.200\\\Rightarrow\dfrac{0.4}{0.4+n_2}=0.200"

"\Rightarrow n_2=1.6\space mol"

Moles of H2O = 1.6 mol

Mass of H2O needed = "1.6\times18=28.8\space g"